# Microbiomes Through the Looking Glass

### v1

@@ Abstract
Bacterial communities are pivotal to maintaining ecological function and preserving the rich tapestry of biological diversity. The rapid development of environmental sequencing technologies, such as metagenomics, has revolutionized our capacity to probe such diversity. However, despite these advances, a theoretical understanding connecting empirical data with ecosystem modelling, in particular in the framework of disordered systems akin to spin glasses, is still in its infancy. Here, we present a comprehensive framework using theories of disordered systems to decode microbiome data, which offers insight into the ecological forces that shape macroecological states. By employing the quenched disordered generalized Lotka-Volterra model, we analyze species abundance data in healthy and diseased human gut microbiomes. Results reveal the emergence of two distinct patterns of species-interaction networks, elucidating the pathways through which dysbiosis may drive microbiome instability. Interaction patterns thus provide a window into the systemic shifts accompanying the transition from health to disease, offering a new perspective on the dynamics of the microbial community. Our findings suggest the potential of disordered systems theory to characterize microbiomes by capturing the essence of ecological interactions and their consequences on stability and functioning, leveraging our understanding of the linkages of dysbiosis and microbial dynamics.
Keywords Dysbiosis  $\cdot$ disordered Lotka-Volterra model  $\cdot$ interaction networks  $\cdot$ bacterial communities  $\cdot$ gastrointestinal diseases
@@ 1 Introduction
Bacterial communities are a fundamental reservoir of ecological functions and biological diversity. They are relevant for any environmental and host-associated ecosystem, ranging from soil to the human gut. In particular, the human gut microbiome, by interacting with host metabolism and immune system Levy et al. (2016); Barroso-Batista et al. (2015); Barreto and Gordo (2023), is a key regulator of human health. Dysbioses are defined as alterations of the healthy microbiome composition associated with the gastrointestinal tract Lloyd-Price et al. (2019); Das and Nair (2019); Nishida et al. (2018). Some recent studies have shown that gut dysbioses are associated with changes in microbial species interaction patterns Bashan et al. (2016); Seppi et al. (2023), while emergent patterns focusing on single species properties Azaele et al. (2016); Grilli (2020), such as species abundance distribution (SAD), fail to characterize altered states of microbiome Seppi et al. (2023); Pasqualini et al. (2023). Microbiomes host various interactions, measured for in-vitro communities Kehe et al. (2021); Venturelli et al. (2018), and are expected to govern natural ones.
Recently, stochastic non-interacting neutral Zeng et al. (2015); Venkataraman et al. (2015); Sala et al. (2016); Sieber et al. (2019) and logistic models Descheemaeker and De Buyl (2020); Grilli (2020); Zaoli and Grilli (2022), have been applied to give a satisfactory description of empirical patterns in different types of communities, ranging from human to soil microbiomes. However, they completely neglect species interactions in the population dynamics, and therefore they do not seem to have the right modelling tools to obtain novel insights from dysbioses models.
@@
For these reasons, several approaches have been proposed in recent years to infer microbial interactions Faust and Raes (2012); Xiao et al. (2017); Camacho-Mateu et al. (2024). However, such inference remains very challenging and problematic in many respects Faisal et al. (2010); Angulo et al. (2017); Tu et al. (2019); Armitage and Jones (2019); Holt (2020): a) the very high dimension of typical microbiome datasets and the lack of longitudinal (long-term) experiments; b) the time-dependent nature of microbial species interactions Pacciani-Mori et al. (2021); c) abiotic factors such as resources, temperature, and pH, can introduce environmental filtering effects Sireci et al. (2022) and induce effective interactions; d) experimental and technical challenges that introduce sampling effects, false positive species Tovo et al. (2020), spurious correlation, and bias in the data Weiss et al. (2016); Gloor et al. (2017); Dohlman and Shen (2019). Therefore, even in a scenario where metagenomic samples are noise- and bias-free, reconstructing species-interaction networks of the underlying microbial dynamics remains a hard task.
@@
An alternative, theoretically more elegant approach was pioneered by May May (1972), proposing to use random matrices to model species interaction networks, i.e., each entry of the adjacency matrix is extracted at random from a given distribution. Given the impossibility of empirically measuring the interaction strengths, the advantage of such an approach is the reduction from $\sim S^{2}$ (if $S$ is the number of species) to just a few parameters (e.g., 2) used to parameterize the distribution Allesina and Tang (2012). In recent years several investigations have applied this framework to the study of population dynamics through generalized Lotka-Volterra (gLV) equations, employing quenched disordered system techniques (also known as "glassy"), mutated from Statistical Physics Bunin (2017); Galla (2018); Biroli et al. (2018); Altieri et al. (2021); Lorenzana and Altieri (2022). The main idea is thus to focus on the ensemble properties of the gLV, which will depend only on a few relevant control parameters (e.g., the mean $\mu$ and the variance $\sigma^{2}$ of the random interactions strengths) Barbier et al. (2018). However, despite a range of very interesting results, this approach has remained confined mainly to purely theoretical domains (but see Hatton et al. (2024)) and has only recently been used to give an interpretation of controlled experiments with microbial communities Hu et al. (2022).
In this work, we therefore provide a proof of concept for the applicability of this framework to human microbiomes. In particular, we will infer the parameters of the quenched dgLV model from healthy and unhealthy cohorts of gut microbiomes. We will show that the different physiological states of the human gut microbiome correspond to different interaction patterns of the dgLV model. We will then introduce quantitative metrics to quantify the role of stability and the contribution of ecological forces to the dynamics.
@@ 2.1 Disordered generalized Lotka-Volterra model
The dgLV model describes the evolution in time of the concentration abundances of a local pool of $S$ interacting species, i.e.,
@@
The replica formalism, a well-known technique in disordered systems, comes into play and allows us to derive a non-interacting Hamiltonian corresponding to the dgLV symmetric interactions $\alpha_{ij}$ (see Methods and the SI, Section B for the complete derivation). In simple words, instead of trying to solve the problem for one random setup, the formalism considers many replicas of the system and averages their behaviors. This approach helps to smooth out the randomness and reveals the typical behavior of the system. We also assume a single equilibrium scenario, known as replica-symmetric (RS) regime. Although the SADs of empirical microbial communities display fat tails, a feature more compatible with the multi-attractors phase of the asymmetric 1RSB Mallmin et al. (2024) case or with the GLV with annealed disorder Suweis et al. (2023), we consider this simplification as it is the regime in which we can obtain explicit analytical relations between the model parameters and the data and where model inversion is feasible.
On the other hand, by using a cavity argument Mezard and Montanari (2009); Altieri and Baity-Jesi (2024), one can analytically derive the SAD of the model (see Methods and SI, Section D):
@@
where the auxiliary variable $\zeta=K-\mu h+\sqrt{q_{0}}\sigma z$ takes the disorder interactions into account; $z$ is a standardized Gaussian variable, $\nu=\beta\lambda>0$, $h=\overline{\langle N\rangle}$, $q_{d}=\overline{\langle N^{2}\rangle}$, $q_{0}=\overline{\langle N\rangle^{2}}$, and $m=1-\beta\sigma^{2}(q_{d}-q_{0})$ that should be greater than zero for the theory to be consistent Biroli et al. (2018).
The RS ansatz can also be characterized by its stability to external perturbations, such as the external temperature, the immigration rate, or the interaction heterogeneity. To investigate the stability of the RS phase, we consider the Hessian matrix of the theory by performing a harmonic expansion of the replicated free energy de Almeida and Thouless (1978). When the leading eigenvalue of the Hessian matrix – suitably diagonalized on the so-called replicon sector – goes below zero, unstable equilibria appear: either the system moves towards a Replica Symmetric Breaking (RSB) phase with multiple locally stable minima, or it develops a marginal full-RSB phase, leading to a hierarchical structure of states and a very complex landscape (also known as glass phase). Therefore, in the latter cases, the RS ansatz no longer represents a thermodynamically stable phase. Based on standard calculations in disordered systems, this leading eigenvalue of the Hessian matrix, called replicon $\mathcal{R}$, can be analytically evaluated as
$\mathcal{R}=\left(\beta\sigma\right)^{2}\left(1-\sigma^{2}\overline{\langle\left(\frac{\partial N}{\partial\xi}\right)^{2}\rangle}\right)=\left(\beta\sigma\right)^{2}\left(1-\left(\beta\sigma\right)^{2}\overline{\left(\langle N^{2}\rangle-\langle N\rangle^{2}\right)^{2}}\right),$ (3)
where $\frac{\partial N}{\partial\xi}$ is the species response to an external perturbation, keeping track of non-extinct species only. For more details, see SI Section C.
@@ 2.2 Data Through the Glass
We now consider cross-sectionally sampled gut microbiomes of two different cohorts: one of healthy and another one of diseased individuals. Here we focus on chronic inflammation syndromes, but, in principle, what we present holds for groups of phenotypically distinct populations. We can idealize the samples of each group as generated from the stationary distribution of the dgLV Eq.(1), with different realizations of disorder $\boldsymbol{\alpha}$, but with shared $\mu$ and $\sigma$ (see Methods). The approach we propose therefore considers all samples in the same group as coming from the same statistical ensemble but, on the other hand, it distinguishes ecological regimes characterized by distinct phenotypes through different statistics (e.g., dissimilar $\mu$ and $\sigma$) of the random species interactions $\boldsymbol{\alpha}$. We aim to test the hypothesis that each cohort will be described by a different set of ecological parameters ($\boldsymbol{\theta}=(\mu,\sigma,T,\lambda)$).
To calculate the order parameters $(h,q_{d},q_{0},K)$ from the data, we thus need to specify how we perform the ensemble and the disorder averages from the data (see Figure 1).
@@
Figure 1: Inference of the dgLV generative model is performed by a moment matching optimization procedure. We aim to infer the free parameters $\boldsymbol{\theta}=(\mu,\sigma,T,\lambda)$ so that to match the order parameters and the average carrying capacity $(h,q_{0},q_{d},K)$. This procedure allows us to extract ecological dynamics information for cross-sectional data of healthy (blue) and diseased (red) microbiomes, which are conceptualized as independent disorder realizations.
@@
Given that we typically lack time series data, and leveraging on an effective mean-field description, we consider estimating the ensemble average as the average among species (that is, different species are realizations of the same underlying stochastic process Azaele et al. (2016), known as neutral hypothesis), i.e., $\langle\ \cdot\ \rangle\sim\frac{1}{S}\sum_{\textit{species}}\cdot$. On the other hand, we assume that the average over the disorder can be computed as a sample average. In other words, within a given phenotype (healthy/unhealthy) each measured microbiome configuration is a sample from the stationary distribution of the dgLV model, with a given realization of the disorder, i.e., ¯$\overline{\ \cdot\ }\sim\frac{1}{R}\sum_{\textit{samples}}\cdot$, where $R$ is the number of samples and typically $R\gg 1$. Moreover, due to our limited knowledge of the fine details governing the interactions in each microbiome, it is reasonable to assume that all communities with a given macrostate – a point in the ($h$, $q_{0}$, $q_{d}$, $K$) space – experience the same demographic noise $T$, immigration rate $\lambda$, and disorder parameters $\mu$ and $\sigma$. Eventually, we can compute the order parameters from the data as:
@@
$\begin{split}h&=\overline{\langle N\rangle}=\frac{1}{R}\sum_{a=1}^{R}\left(\frac{1}{S}\sum_{j=1}^{S}N_{j,a}\right)\ ,\\ q_{d}&=\overline{\langle N^{2}\rangle}=\frac{1}{R}\sum_{a=1}^{R}\left(\frac{1}{S}\sum_{j=1}^{S}N^{2}_{j,a}\right)\ ,\\ q_{0}&=\overline{\langle N\rangle^{2}}=\frac{1}{R}\sum_{a=1}^{R}\left(\frac{1}{S}\sum_{j=1}^{S}N_{j,a}\right)^{2}\,\end{split}$ (4)
where $N_{a,j}$ gives the population density of species $j$ in sample $a$.
@@
To evaluate the carrying capacities $K_{i}$ from the data, as done in most of the works using dGLV Bunin (2017); Biroli et al. (2018); Altieri et al. (2021); Suweis et al. (2023); Mallmin et al. (2024), we assume that in each cohort, all species are characterized by the same carrying capacity $K_{i}=K$. Here, we assign the value $K$ as the average maximum relative abundance of species among the available samples for each cohort, thus $K=\frac{1}{S}\sum_{j=1}^{S}\max_{a}N_{j,a}$.
Note also that, from metagenomics, we only have access to compositional data for species abundances Pasqualini et al. (2023). This is crucial to properly treat different samples and end up with a consistent analysis. Moreover, this also allows us to give an ecological interpretation of some of the order parameters. In particular, thanks to the compositionality of the data, we have that $h=\overline{S^{-1}}$ and $q_{0}=\overline{S^{-2}}$ (where $\overline{S}$ is known as the average local diversity), while $q_{d}$ provides information on pairwise products between species abundances within each replica. Figure 2 shows the values of such order parameters between healthy and unhealthy cohorts with respect to two randomized cohorts. In the data, we observe a systematic difference between healthy and unhealthy cohorts, pointing to a higher average local diversity in healthy samples (panels A, B), as also observed in Pasqualini et al. (2023). Panel C highlights the higher $q_{d}$ in unhealthy patients, a signature of the weakening of the species interactions in those samples. We will investigate further this aspect in the following section, by inferring the species interactions in both cohorts.
From the modelling point of view, it can be rigorously shown that constraining the abundances to be normalized to one corresponds to introducing an additional Lagrange multiplier in the expression of the free energy. However, this only acts on the linear term contributing to shifting the mean of the random interactions, $\mu$, but does not have any influence on the heterogeneity $\sigma$, therefore on the phase diagram. For more details, we refer the interested reader to the SI (Section B1) along with Altieri et al. (2021). In Altieri et al. (2021) the authors also showed how a global normalizing constraint on species abundances reflects in a one-to-one mapping with the random replicant model Biscari and Parisi (1995) [^1] The replicator equations originally introduced by Diederich and Opper (1989) and recast within the replica formalism Biscari and Parisi (1995); Altieri et al. (2021) describe the dynamics of an ensemble of replicants that evolve via random couplings..
@@
Figure 2: Order parameters inferred from the data using Eqs. (4). Panel A) $h$ B) $q_{0}$ and C) $q_{d}$ in healthy (blue) and unhealthy (red) cohorts. The gray points show the values for the corresponding null models where sample labels have been randomized. Standard deviations have been computed over 5000 realizations.

### v2

@@ Abstract
The rapid advancement of environmental sequencing technologies, such as metagenomics, has significantly enhanced our ability to study microbial communities. The eubiotic composition of these communities is crucial for maintaining ecological functions and host health. Species diversity is only one facet of a healthy community’s organization; together with abundance distributions and interaction structures, it shapes reproducible macroecological states, i.e., joint statistical fingerprints that summarize whole-community behavior. Despite recent developments, a theoretical framework connecting empirical data with ecosystem modeling is still in its infancy, particularly in the context of disordered systems. Here, we present a novel framework that couples statistical-physics tools for disordered systems with metagenomic data, explicitly linking diversity, interactions, and stability to define and compare these macroecological states. By employing the generalized Lotka-Volterra model with random interactions, we reveal two different emergent patterns of species-interaction networks and species abundance distributions for healthy and diseased microbiomes. On the one hand, healthy microbiomes have similar community structures across individuals, characterized by strong species interactions and abundance diversity consistent with neutral stochastic fluctuations. On the other hand, diseased microbiomes show greater variability driven by deterministic factors, thus resulting in less ecologically stable and more divergent communities. Our findings suggest the potential of disordered system theory to characterize microbiomes and to capture the role of ecological interactions on stability and functioning.
@@ 1 Introduction
Microbial communities are a fundamental reservoir of ecological functions and biological diversity. They are relevant for any environmental and host-associated ecosystem, ranging from soil to the human gut. In particular, the human gut microbiome, by interacting with the host’s metabolism and immune system Levy et al. (2016); Barroso-Batista et al. (2015); Barreto and Gordo (2023), is a key regulator of human health. Dysbiosis is defined as an alteration in the composition of healthy microbiomes associated with the gastrointestinal tract Lloyd-Price et al. (2019); Das and Nair (2019); Nishida et al. (2018).
Recently, stochastic non-interacting neutral Zeng et al. (2015); Venkataraman et al. (2015); Sala et al. (2016); Sieber et al. (2019)Descheemaeker and De Buyl (2020); Grilli (2020); Zaoli and Grilli (2022), and logistic models have been successfully used to describe empirical patterns such as species abundance distribution (SAD) and species presence/absence statistics in different types of communities. However, focusing on single-species properties Azaele et al. (2016); Grilli (2020) fails to characterize altered states of microbiome Seppi et al. (2023); Pasqualini et al. (2024). In particular, recent studies indicate that gut dysbiosis is associated with shifts in microbial species interaction patterns. Network-level properties of species interactions – such as the balance of positive and negative interactions, average interaction strength, and connectivity of the inferred interaction matrix – have been shown to systematically differ between healthy and dysbiotic gut communities (see, for instance, Bashan et al. (2016); Seppi et al. (2023)). Pairwise species interactions have been quantified in simplified in-vitro consortia Kehe et al. (2021); Venturelli et al. (2018); we assume that the same classes of interactions also operate – albeit in a more complex form – in the native gut microbiome.
@@
For these reasons, several approaches have been proposed in recent years to infer microbial interactions Faust and Raes (2012); Xiao et al. (2017); Camacho-Mateu et al. (2024). However, such inference protocols remain very challenging and problematic in many respects Faisal et al. (2010); Angulo et al. (2017); Tu et al. (2019); Armitage and Jones (2019); Holt (2020): a) the very high dimension of typical microbiome datasets and the lack of longitudinal (long-term) experiments; b) the time-dependent nature of microbial species interactions Pacciani-Mori et al. (2021); c) abiotic factors such as resources, temperature, and pH, can introduce environmental filtering effects Sireci et al. (2022) and induce effective interactions; d) experimental and technical challenges that introduce sampling effects, false positive species Tovo et al. (2020), spurious correlation, and bias in the data Weiss et al. (2016); Gloor et al. (2017); Dohlman and Shen (2019). Therefore, even in a scenario where metagenomic samples are noise- and bias-free, reconstructing species-interaction networks of the underlying microbial dynamics remains a hard task.
An alternative, theoretically more elegant approach was pioneered by May May (1972), proposing to use random matrices to model species interaction networks, i.e., each entry of the adjacency matrix is extracted at random from a given distribution. Given the impossibility of empirically measuring the interaction strengths, the advantage of such an approach is the reduction from $\sim S^{2}$ (if $S$ is the number of species) to just a few parameters (e.g., 2) used to parameterize the distribution Allesina and Tang (2012).
1.1 Insights from Disordered Systems Theory
Although previous studies have largely focused on neutral theory and logistic models – including variants like the Stochastic Logistic Model that account for environmental fluctuations – they often neglect inter-species interactions, limiting their ability to reproduce macroecological patterns at a global scale. Moreover, higher-order correlations among species are known to generate non-trivial effects, such as the emergence of persistent fluctuations or chaotic dynamics due to the non-reciprocity of interactions.
@@ 1.1 Insights from Disordered Systems Theory
In recent years, increasing attention has been devoted to studying population dynamics through the generalized Lotka-Volterra (gLV) equations, employing disordered systems techniques (also known as glassy), such as replica and cavity methods and Dynamical Mean-Field Theory approaches, originally developed in the context of statistical physics Bunin (2017); Galla (2018); Biroli et al. (2018); Altieri et al. (2021); Lorenzana and Altieri (2022). Indeed, due to the inherently high dimensionality of microbial datasets, random matrix theory and methods from disordered systems turn out to be particularly well suited. A striking feature of their application to ecological dynamics is that the resulting properties and dynamical regimes do not depend on species-specific details, provided that species are statistically equivalent under relabeling or time-averaging. Phase diagrams can thus be established in terms of a few effective parameters (e.g., the mean $\mu$ and the variance $\sigma^{2}$ of the random interactions strengths) Barbier et al. (2018). However, despite a range of very interesting results, this approach has remained confined mainly to purely theoretical domains (but see Hatton et al. (2024)) and has only recently been used to give an interpretation of controlled experiments with synthetic microbial communities Hu et al. (2022).
@@ 2 Results
We now provide a proof of concept of the applicability of a high-dimensional disordered setting to human microbiomes. Specifically in the following section, we introduce the disordered generalized Lokta-Volterra (dgLV) model, infer its parameters from healthy and unhealthy cohorts of gut microbiomes, and test whether the resulting interaction patterns and stability metrics discriminate healthy from diseased macro-ecological states. Finally, we will introduce quantitative metrics to define microbiome stability and estimate the contribution of distinct ecological forces to the dynamics.
@@ 2.1 Disordered generalized Lotka-Volterra model
The dgLV model describes the time evolution of the concentration abundances of a local pool of $S$ interacting species, i.e.,
@@
where $N_{i}$ is the populations of species $i$-th, $K_{i}$ is its carrying capacity, and $r_{i}$ the growth rate where $\rho_{i}=\frac{r_{i}}{K_{i}}$ is a constant, which we will assume will not depend on the species, i.e. $\rho_{i}=\rho$. The coefficients $\alpha_{ij}$ are random i.i.d random variables with $\mathbb{E}[\alpha_{ij}]=\mu/S$ and $\mathbb{E}[\alpha_{ij}]^{2}-\mathbb{E}^{2}[\alpha_{ij}]=\sigma^{2}/S$. We also include an immigration species-independent rate $\lambda$, which will be treated as a reflecting wall mathematically regularizing the problem, and a demographic noise term Altieri et al. (2021) $\eta_{i}$, with $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=2N_{i}T\delta_{ij}\delta(t-t^{\prime})$, and where $\delta$ denotes the Dirac delta and $T$ sets the scale of noise intensity. For notational purposes, we also introduce its inverse $\beta=T^{-1}$. Finally, we require that the interactions are symmetric, i.e., $\alpha_{ij}=\alpha_{ji}$. We can thus map this problem into an equilibrium thermodynamic one that is exactly solvable. As shown in the Appendix, section S1, it is possible to justify this symmetric assumption and a linear dependence between the growth rate and the carrying capacity, namely $r_{i}=\rho K_{i}$, by considering the quasi-stationary approximation of the MacArthur consumer-resource model. In the following, we set $\rho=1$ (following also Biroli et al. (2018)).
@@
The replica formalism, a well-known technique in disordered systems, comes into play and allows us to derive a non-interacting Hamiltonian corresponding to the dgLV symmetric interactions $\alpha_{ij}$ (see Methods and Appendix, Section S2, for a complete derivation). In simple words, instead of trying to solve the problem for one random setup, the formalism considers many replicas of the system and averages their behaviors. This approach helps to smooth out the randomness and reveals the typical behavior of the system. We also assume a single equilibrium scenario, known as replica-symmetric (RS) regime. Although the SADs of empirical microbial communities display fat tails, a feature more compatible with the multi-attractors phase of the asymmetric 1RSB Mallmin et al. (2024) case or with the gLV with time-dependent disordered interactions Suweis et al. (2024) (annealed version), we consider this simplification as it is the regime in which we can obtain explicit analytical relations between the model parameters and the data and where model inversion is feasible. By employing a cavity argument Mezard and Montanari (2009); Altieri and Baity-Jesi (2024), one can indeed analytically derive the SAD of the model (see Methods and Appendix, Section S4):
@@
where the auxiliary variable $\zeta=K-\mu h+\sqrt{q_{0}}\sigma z$ takes the disorder interactions into account; $z$ is a standardized Gaussian variable, $\nu=\beta\lambda>0$. We introduce, respectively, the mean abundance $h=\overline{\langle N\rangle}$, the self-overlap $q_{d}=\overline{\langle N^{2}\rangle}$, and the overlap $q_{0}=\overline{\langle N\rangle^{2}}$, which we will collectively refer to as order parameters in the following sections. Specifically, $q_{0}$ measures the similarity between two different configurations at stationarity of the system with two different disorder realizations, while $q_{d}$ measures the similarity of two stationary configurations generated with the same disorder realization. Finally, we introduce a constant that, inspired by the Field Theory jargon, we dub as mass $m=1-\beta\sigma^{2}(q_{d}-q_{0})>0$ for the theory to be consistent Biroli et al. (2018).
The RS ansatz can also be characterized by its stability to external perturbations, such as the external temperature, the immigration rate, or the interaction heterogeneity. To investigate the stability of the RS phase, we consider the Hessian matrix of the theory by performing a harmonic expansion of the replicated free energy, as originally pointed out in de Almeida and Thouless (1978). When the leading eigenvalue of the Hessian matrix, the so-called replicon mode, goes below zero, unstable equilibria appear: either the system moves towards a One-step Replica Symmetric Breaking (1-RSB) phase with multiple locally stable minima, or it develops a marginal full-RSB phase, leading to a hierarchical structure of states and therefore to an extremely rough landscape (also known as Gardner amorphous-like phase). In other words, the RS ansatz no longer represents a thermodynamically stable phase in the latter cases.
Based on standard calculations in disordered systems, the replicon mode $\mathcal{R}$ can be analytically evaluated and reads [^1] The replicon eigenvalue refers to a particular type of fluctuation around the saddle-point (mean-field) solution in the replica framework. When diagonalizing the Hessian matrix of the replicated free energy, fluctuations split into three sectors: longitudinal, anomalous, and replicon. The replicon mode is the most sensitive to criticality signaling – by its vanishing trend – the emergence of many nearly-degenerate states. It essentially describes how ‘soft’ the system is to microscopic rearrangements in configuration space.:
$\mathcal{R}=\left(\beta\sigma\right)^{2}\left(1-\sigma^{2}\overline{\langle\left(\frac{\partial N}{\partial\xi}\right)^{2}\rangle}\right)=\left(\beta\sigma\right)^{2}\left(1-\left(\beta\sigma\right)^{2}\overline{\left(\langle N^{2}\rangle-\langle N\rangle^{2}\right)^{2}}\right),\$(3)
where $\frac{\partial N}{\partial\xi}$ is the species response to an external perturbation, keeping track of non-extinct species only. For more details, see S3 in the Appendix.
@@ 2.2 Data Through the Glass
We now consider cross-sectionally sampled gut microbiomes of two different cohorts: one of healthy and another of diseased individuals. Here we focus on chronic inflammation syndromes, but, in principle, what we present holds for groups of phenotypically distinct populations. We can idealize the samples of each group as generated from the stationary distribution of the dgLV Eq. (1), with different realizations of the disorder $\boldsymbol{\alpha}$, but with shared $\mu$ and $\sigma$ (see Material and Methods). The approach we propose considers all samples in the same group as coming from the same statistical ensemble. On the other hand, it distinguishes ecological regimes characterized by distinct phenotypes through different statistics (e.g., dissimilar $\mu$ and $\sigma$) of the random species interactions $\boldsymbol{\alpha}$. We aim to test the hypothesis that each cohort will be described by a different set of ecological parameters, precisely with $\boldsymbol{\theta}=(\mu,\sigma,T,\lambda)$.
@@
Figure 1: The inference protocol of the dgLV generative model is performed by a moment matching optimization procedure. We aim to infer the free parameters $\boldsymbol{\theta}=(\mu,\sigma,T,\lambda)$ – shown on the right – so that to match the mean abundance, higher-order correlations between species abundances, and the average carrying capacity for the two cohorts, i.e. $(h,q_{0},q_{d},K)$ – on the left. This procedure allows us to extract ecological dynamics information for cross-sectional data of healthy (blue) and diseased (red) microbiomes, which are conceptualized as independent disorder realizations.
@@
To calculate the order parameters $(h,q_{d},q_{0},K)$ from the data, we thus need to specify how we perform the ensemble and the disorder averages empirically (see Figure 1). Since time series data are rarely available, we rely on an effective mean-field description and estimate the ensemble average by averaging over species (that is, different species are realizations of the same underlying stochastic process Azaele et al. (2016), known as neutral hypothesis), i.e., $\langle\ \cdot\ \rangle\sim\frac{1}{S}\sum_{\textit{species}}\cdot$. Then, we assume that the average over the disorder can be computed as a sample average. In other words, within a given phenotype (healthy/unhealthy), each measured microbiome configuration is a sample from the stationary distribution of the dgLV model, with a given realization of the disorder, i.e., $\overline{\ \cdot\ }\sim\frac{1}{R}\sum_{\textit{samples}}\cdot$, where $R$ is the number of samples and typically $R\gg 1$. Moreover, due to our limited knowledge of the fine details governing the interactions in each microbiome, it is reasonable to assume that all communities with a given macrostate – a point in the ($h$, $q_{0}$, $q_{d}$, $K$) space – experience the same demographic noise $T$, immigration rate $\lambda$, and disorder parameters $\mu$ and $\sigma$. Eventually, we can compute the order parameters from the data as:
@@
$\begin{split}h&=\overline{\langle N\rangle}=\frac{1}{R}\sum_{a=1}^{R}\left(\frac{1}{S_{a}}\sum_{j=1}^{S}N_{j,a}\right)\ ,\\ q_{d}&=\overline{\langle N^{2}\rangle}=\frac{1}{R}\sum_{a=1}^{R}\left(\frac{1}{S_{a}}\sum_{j=1}^{S}N^{2}_{j,a}\right)\ ,\\ q_{0}&=\overline{\langle N\rangle^{2}}=\frac{1}{R}\sum_{a=1}^{R}\left(\frac{1}{S_{a}}\sum_{j=1}^{S}N_{j,a}\right)^{2}\,\end{split}$ (4)
where $N_{j,a}$ represents the population density of species $j$ in sample $a$.
@@
To evaluate the carrying capacities $K_{i}$ from the data, as done in most of the works using dgLV equations Bunin (2017); Biroli et al. (2018); Altieri et al. (2021); Suweis et al. (2024); Mallmin et al. (2024), we assume that in each cohort all species are characterized by the same carrying capacity $K_{i}=K$. Here, we define $K$ as the average of the maximum relative abundances of species across the available samples for each cohort, so that $K=\frac{1}{S}\sum_{j=1}^{S}\max_{a}N_{j,a}$.
@@
Figure 2: Order parameters inferred from the data using Eqs. (4). Panel a) shows $h$, b) $q_{d}$, and panel c) $q_{0}$ in healthy (blue) and unhealthy (red) cohorts. The gray points show the values for the corresponding null models where sample labels have been randomized. Standard deviations have been computed over 5000 realizations.
Note also that, from metagenomics, we only have access to compositional data for species abundances Pasqualini et al. (2024). This is crucial to properly treat different samples and end up with a consistent analysis. Moreover, this approach allows us to give an ecological interpretation of some of the order parameters. In particular, because of the compositionality of the data, we have that $h=\overline{S^{-1}}$ and $q_{0}=\overline{S^{-2}}$, while $q_{d}$ provides information on pairwise products between species abundances within each replica. Figure 2 shows the values of such order parameters between healthy and unhealthy cohorts for two randomized instances. In the data, we observe a systematic difference between healthy and unhealthy cohorts, pointing to a higher average local diversity in healthy samples (panels a, c), as also observed in Pasqualini et al. (2024). Panel b highlights a higher $q_{d}$ in unhealthy patients, a signature of the weakening species interactions in those samples. We will further investigate this aspect by inferring the species interactions in both cohorts.
From the modeling point of view, it can be rigorously shown that constraining the abundances to be normalized to one corresponds to introducing a Lagrange multiplier in the expression of the free energy. However, this only acts on the linear term, contributing to shifting the mean of the random interactions, $\mu$, but does not affect the heterogeneity $\sigma$, and therefore the phase diagram overall. For more details, we refer the interested reader to the Appendix (end of Section S2) along with Altieri et al. (2021). In Altieri et al. (2021) the authors also showed how a global normalizing constraint on species abundances reflects in a one-to-one mapping with the random replicant model Biscari and Parisi (1995) [^2] The replicator equations originally introduced by Diederich and Opper (1989) and recast within the replica formalism Biscari and Parisi (1995); Altieri et al. (2021) describe the dynamics of an ensemble of replicants that evolve via random couplings..